OpenStudy (anonymous):

log base 32 of (1/4) log base 32 of (1/4) @Mathematics

OpenStudy (anonymous):

your job is to solve \[32^{x}=\frac{1}{4}\] do you know it? i can show you how to cheat with a calculator if you like

OpenStudy (anonymous):

no I don't know it. Please show me how to do it on a calculator

myininaya (myininaya):

\[\log_{32}(\frac{1}{4})=x\] satellite wrote this an exponential form

OpenStudy (anonymous):

first how you do it without a calculator. \[32=2^5\] and \[\frac{1}{4}=2^{-2}\] so you can solve \[5x=-2\] for x

myininaya (myininaya):

\[x=\frac{\ln(\frac{1}{4})}{\ln(32)}=\frac{\ln(1)-\ln(4)}{\ln(32)}=\frac{0-\ln(4)}{\ln(32)}=-\frac{\ln(2^2)}{\ln(2^5)}\] \[=-\frac{2 \cdot \ln(2)}{5 \cdot \ln(2)}=\frac{-2}{5}\]

myininaya (myininaya):

i like satellite's way better

myininaya (myininaya):

i used change of base formula

myininaya (myininaya):

\[\log_b(x)=\frac{\ln(x)}{\ln(b)}\]

OpenStudy (anonymous):

if you are going to use a calculator then just plug in myininaya's first line and type in \[\ln(\frac{1}{4})\div\ln(32)\]

OpenStudy (anonymous):

or even \[\ln(.25)\div\ln(32)\] or \[\log(,25)\div\log(32)\] but you are supposed to probably write my first answer on a test

OpenStudy (anonymous):

I've never learned the change base formula so I'll probably stick with the first method. And thanks for the calculator method

OpenStudy (anonymous):

yw

myininaya (myininaya):

thats interesting i wonder how your teacher expects you to enter it into the calculator without knowing the change of base formula

OpenStudy (anonymous):

that was my suggestion. i am fairly sure that the teacher wants jennifer to solve \[32^x=\frac{1}{4}\]

myininaya (myininaya):

ok

OpenStudy (anonymous):

it is true. She doesn't want us to use calculators most of the time